# Bloch-Beilinson conjectures for Hecke characters and Eisenstein cohomology of Picard surfaces

### v2

@@ 2.8. Local systems on $\mathrm{S}_{K}({\mathbb{C}})$
Fix one of the locally symmetric spaces $S_{\Gamma_{i}}$ arising as in (20), and call it $\mathrm{S}_{\Gamma}$. Its fundamental group is identified to $\Gamma$. Hence, any finite dimensional representation of $\mathrm{G}$, by restriction to $\Gamma$, naturally defines a local system on $\mathrm{S}_{\Gamma}$. For any irreducible representation $V_{\lambda}$ of $\mathrm{G}$ of highest weight $\lambda$, we denote by the same symbol the resulting local system. One gets an isomorphism
(21) $H^{\bullet}(\Gamma,V_{\lambda})\cong H^{\bullet}(\mathrm{S}_{\Gamma},V_{\lambda}),$
@@
We will denote by $\mathop{\rm MHS}\nolimits$ the category of mixed ${\mathbb{Q}}$-Hodge structures. Moreover, we will need to work with a generalization of this category: for ${\mathbb{K}}$ a finite extension of ${\mathbb{Q}}$, the category $\mathop{\rm MHS}\nolimits_{{\mathbb{K}}}$ of mixed Hodge structures with coefficients in ${\mathbb{K}}$, or of mixed ${\mathbb{K}}$-Hodge structures, is the pseudo-abelian completion of the category whose objects are the same as $\mathop{\rm MHS}\nolimits$ and where the morphisms between objects $A,B$ are defined by $\mathop{\rm Hom}\nolimits_{\mathop{\rm MHS}\nolimits}(A,B)\otimes{\mathbb{K}}$. When we do not want to specify the number field ${\mathbb{K}}$, we will speak of Hodge structures with $\overline{{\mathbb{Q}}}$-coefficients. We leave to the reader the task of guessing and working out the precise definition of the category $\mathop{\rm MHS}\nolimits_{\overline{{\mathbb{Q}}}}$. There are obvious notions of weights and types for a $\overline{{\mathbb{Q}}}$-Hodge structure. Notice that the types of a pure $\overline{{\mathbb{Q}}}$-Hodge structure do not necessarily satisfy Hodge symmetry.
@@
(23) $w=r$
This property is the reason for our choice of parametrizing characters as in (12). In particular, $V_{k}$ is endowed with a pure $F$-Hodge structure of weight $k$.
@@
By general mixed Hodge theory, the cohomology groups $H^{\bullet}(\mathrm{S}_{K},V_{\lambda})$ are endowed with a canonical mixed $F$-Hodge structure.
@@ 3.1. Algebraic Hecke characters
An algebraic Hecke character on a ${\mathbb{Q}}$-torus $\mathcal{T}$ is a continuous homomorphism
@@
We will often refer to a Hecke character on $\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}{\mathbb{G}}_{m,F}$ as a Hecke character of $F$.
A Hecke character on $\mathcal{T}$ is algebraic if there exist a character
@@
××$\begin{array}[]{ccc}{\mathbb{C}}^{\times}&\rightarrow&{\mathbb{C}}^{\times}\\ z&\mapsto&z^{-a}\bar{z}^{-b}\end{array}$
In this case, we say that $(a,b)$ is the infinity type of $\phi$, and $w(\phi):=a+b$ is its weight. Note that the finite part $\phi_{f}$ of an algebraic Hecke character takes values in $\overline{{\mathbb{Q}}}^{\times}$, see e.g. [16, Sec. 2.5]. More precisely, it takes values in a number field $F^{\prime}$, called its field of values.
In our case, algebraic Hecke characters on $\mathrm{T}^{\prime}$ will be identified to Hecke characters of $F$ through the isomorphism (10). As at the end of subsection 2.4, consider the natural embedding $\mathrm{T}^{\prime}\rightarrow\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}\mathrm{T}^{\prime}_{F}$. Fix a character $\chi:\mathrm{T}^{\prime}_{F}\rightarrow{\mathbb{G}}_{m,F}$, and write it $\chi=(\kappa_{1},\kappa_{2})$ in the parametrization of subsection 2.3. Then $\chi$ restricts, under the previous embedding, to the map given on ${\mathbb{R}}$-points by
$\left(\begin{array}[]{ccc}\bar{z}&&\\ &\bar{z}^{-1}z&\\ &&z^{-1}\end{array}\right)\mapsto z^{\kappa_{2}}\bar{z}^{\kappa_{1}-\kappa_{2}}\,,$
With these conventions, the weight of an Hecke character $\phi$ of $F$ of type $\chi=(\kappa_{1},\kappa_{2})$ is
@@
For later use, let us recall some basic properties of the $L$-function $L(\phi,s)$ of an algebraic Hecke character $\phi$ of $F$. The region in which $L(\phi,s)$ is expressed as an absolutely convergent Euler product is the half-plane $\mbox{Re}\ s>1+\frac{w(\phi)}{2}$. In this region, the $L$-function is therefore holomorphic, and moreover known to be non-zero. Multiplying the Euler product by appropriate $\Gamma$-factors, we obtain a completed L-function $\Lambda(\phi,s)$ which admits a meromorphic continuation to the whole of ${\mathbb{C}}$. The function $\Lambda(\phi,s)$ and the meromorphically continued $L(\phi,s)$ are entire unless $\phi(\cdot)=|\cdot|_{{\mathbb{I}}_{F}}^{t}$ for some $t\in{\mathbb{C}}$; in the latter case, $\Lambda(\phi,s)$ has first-order poles precisely in $s=-t$ and $s=1-t$, and $L(\phi,s)$ has a unique first-order pole in $s=1-t$. Finally, there exists a holomorphic function $\epsilon(s)$ such that $\Lambda(\phi,s)$ satisfies the functional equation
@@ 3.2. Motives for Hecke characters
By [13, Prop. 3.5], there exists a Chow motive $M_{\phi}$ with coefficients in a finite extension of $F^{\prime}$, whose realizations are the Hodge structure $H_{\phi}$ and the compatible system $(M_{\phi,\ell})_{\ell}$. One then has $L(\phi,s)=L(M_{\phi},s)$. The absolute Hodge motive corresponding to $M_{\phi}$ is the one attached to $\phi$ in [32, Ch.1, Thm. 4.1]. Recall that by [13], for every abelian variety $A$, there exists a canonical Chow motive $h^{i}(A)$ realising to the $i$-th cohomology of $A$ in any Weil cohomology theory. Then, the motives $M_{\phi}$ have the following form: if the weight $w$ of $\phi$ is positive, there exist an abelian variety $A_{\phi}$ over $F$, with complex multiplication by a commutative semisimple algebra $T$ over $F$, and an idempotent element $e_{\phi}$ of $T^{\otimes w}$, such that
$M_{\phi}=(h^{1}(A_{\phi})^{\otimes w},e_{\phi})=(h^{w}(A^{w}_{\phi}),e_{\phi})$
Moreover, there exists an isomorphism ([22, Thm. 3.3.1])
$h^{w}(A_{\phi}^{w})\simeq\mathop{\rm Sym}\nolimits^{w}h^{1}(A_{\phi}^{w})$
Since
$\mathop{\rm Sym}\nolimits^{w}h^{1}(A_{\phi}^{w})\simeq\mathop{\rm Sym}\nolimits^{w}h^{1}(A_{\phi})^{\oplus w}$
one gets a projector $e_{1,w}$ defining $h^{1}(A_{\phi})^{\otimes w}$ as a subobject of $h^{w}(A_{\phi}^{w})$. Hence, one can rewrite
$M_{\phi}=(h^{w}(A_{\phi}^{w}),e_{\phi}e_{1,w})$
For any embedding $F\hookrightarrow{\mathbb{C}}$, we can consider the singular cohomology spaces $H^{*}(A^{w}_{\phi})$, with coefficients in a suitable number field, of the analytification of $A^{w}_{\phi}$ with respect to the chosen embedding. This space has a canonical pure Hodge structure of weight $w$, with coefficients in the same number field. The Hodge structure $H_{\phi}$ can be then described as
(25) $H_{\phi}=e_{\phi}e_{1,w}H^{w}(A_{\phi}^{w})$

### v4

@@ 2.8. Local systems on $\mathrm{S}_{K}({\mathbb{C}})$
Fix one of the locally symmetric spaces $S_{\Gamma_{i}}$ arising as in (21), and call it $\mathrm{S}_{\Gamma}$. Its fundamental group is identified with $\Gamma$. Hence, any finite dimensional representation of $\mathrm{G}$, by restriction to $\Gamma$, naturally defines a local system on $\mathrm{S}_{\Gamma}$. For any irreducible representation $V_{\lambda}$ of $\mathrm{G}$ of highest weight $\lambda$, we denote by the same symbol the resulting local system. One gets an isomorphism
(22) $H^{\bullet}(\Gamma,V_{\lambda})\cong H^{\bullet}(\mathrm{S}_{\Gamma},V_{\lambda})\,.$
@@
For the notions and facts that we will use about pure and mixed Hodge structures and variations of Hodge structures, we refer the reader to [33]. Here, we mostly set up our notations.
We will denote by $\mathop{\rm MHS}\nolimits$ the category of mixed $\mathbb{Q}$}-Hodge structures. Moreover, we will need to work with a generalization of this category: for $\mathbb{K}$} a finite extension of $\mathbb{Q}$}, the category $\mathop{\rm MHS}\nolimits_{{\mathbb{K}}}$ of mixed Hodge structures with coefficients in $\mathbb{K}$}, or of mixed $\mathbb{K}$}-Hodge structures, is the pseudo-abelian completion of the category whose objects are the same as $\mathop{\rm MHS}\nolimits$ and where the morphisms between objects $A,B$ are defined by $\mathop{\rm Hom}\nolimits_{\mathop{\rm MHS}\nolimits}(A,B)\otimes{\mathbb{K}}$. Letting the number field $\mathbb{K}$} vary, one obtains a category of Hodge structures with $\overline{{\mathbb{Q}}}$-coefficients. There are obvious notions of weights and types for a $\overline{{\mathbb{Q}}}$-Hodge structure. Notice that the types of a pure $\overline{{\mathbb{Q}}}$-Hodge structure do not necessarily satisfy Hodge symmetry.
@@
(24) w=r.w=r\,.
This property is the reason we chose to parametrize characters as in (13). In particular, $V_{k}$ is endowed with a pure $F$-Hodge structure of weight $k$.
@@
There exists a sheafified version of mixed Hodge theory, including the theory of variations of Hodge structures as a special case. The sheaf-theoretic category that one considers is then called the category of mixed Hodge modules. They are equipped with a six functor formalism (in the sense studied in [18]; for example, one can pullback and pushforward mixed Hodge modules along arbitrary morphisms). We will only need to use this formalism as a black box. In particular, it implies that the cohomology groups $H^{\bullet}(\mathrm{S}_{K},V_{\lambda})$ are endowed with a canonical mixed $F$-Hodge structure.
@@ 3.1. Algebraic Hecke characters
A Hecke character on a $\mathbb{Q}$}-torus $\mathcal{T}$ is a continuous homomorphism
@@
We will often refer to a Hecke character on $\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}{\mathbb{G}}_{m,F}$ as a Hecke character of $F$, which in this case is simply a continuous homomorphism
$F^{\times}\backslash\mathbb{I}_{F}\rightarrow{\mathbb{C}}^{\times}.$
A Hecke character $\phi$ on $\mathcal{T}$ is algebraic if there exists a character
@@
$\begin{array}[]{ccc}{\mathbb{C}}^{\times}&\rightarrow&{\mathbb{C}}^{\times}\\ z&\mapsto&z^{-a}\bar{z}^{-b}\end{array}$
In this case, we say that $(a,b)$ is the infinity type of $\phi$, and $w(\phi):=a+b$ is its weight.
Remark 3.1.
The notion of weight can be defined for Hecke characters over any torus, in such a way that a Hecke character $\phi$ on a torus $\mathcal{T}$ is of weight $w(\phi)$ if and only if, denoting by $\bar{\phi}$ the Hecke character obtained by composing $\phi$ with complex conjugation in $\mathbb{C}$},
(25) $\phi\cdot\bar{\phi}=|\cdot|_{{\mathbb{I}}_{F}}^{-w(\phi)}.$
Now suppose that $\phi$ is a Hecke character of $F$ and denote by $\phi_{{\mathbb{Q}}}$ its restriction to $\mathbb{I}$$_{{\mathbb{Q}}}$. From the above discussion and from the fact that the restriction to $\mathbb{R}$$^{\times}$ of $|\cdot|_{{\mathbb{I}}_{F},\infty}$ is equal to the square of $|\cdot|_{{\mathbb{I}}_{\mathbb{Q}},\infty}$, it is clear that
$w(\phi_{{\mathbb{Q}}})=2w(\phi).$
Note that the finite part $\phi_{f}$ of an algebraic Hecke character takes values in $\overline{{\mathbb{Q}}}^{\times}$, see, e.g. [22, Sect. 2.5]. More precisely, it takes values in a number field $F^{\prime}$, called its field of values.
Recall the maximal torus $\mathrm{T}^{\prime}$ of $\mathrm{G}^{\mathop{\rm\tiny{der}}\nolimits}$, as introduced in Subsection 2.2. We will mostly work with algebraic Hecke characters on $\mathrm{T}^{\prime}$. They will be identified with Hecke characters of $F$ through the isomorphism (11). As at the end of Subsection 2.4, consider the natural embedding $\mathrm{T}^{\prime}\rightarrow\mathop{\rm Res}\nolimits_{F|{\mathbb{Q}}}\mathrm{T}^{\prime}_{F}$. Fix a character $\chi:\mathrm{T}^{\prime}_{F}\rightarrow{\mathbb{G}}_{m,F}$, and write it $\chi=(\kappa_{1},\kappa_{2})$ in the parameterization of Subsection 2.3. Then $\chi$ restricts, under the previous embedding, to the map given on $\mathbb{R}$}-points by
$\left(\begin{array}[]{ccc}\bar{z}&&\\ &\bar{z}^{-1}z&\\ &&z^{-1}\end{array}\right)\mapsto z^{\kappa_{2}}\bar{z}^{\kappa_{1}-\kappa_{2}}\,.$
With these conventions, the weight of a Hecke character $\phi$ of $F$ of type $\chi=(\kappa_{1},\kappa_{2})$ is
@@
For later use, let us recall some basic properties of the $L$-function $L(\phi,s)$ of an algebraic Hecke character $\phi$ of $F$ (see, e.g., [39, Ch. 0, Sect. 6]). The region in which $L(\phi,s)$ is expressed as an absolutely convergent Euler product is the half-plane $\mbox{Re}\ s>1+\frac{w(\phi)}{2}$. In this region, the $L$-function is therefore holomorphic, and moreover known to be non-zero. Multiplying the Euler product by appropriate $\Gamma$-factors, we obtain a completed L-function $\Lambda(\phi,s)$ which admits a meromorphic continuation to the whole of $\mathbb{C}$}. The function $\Lambda(\phi,s)$ and the meromorphically continued $L(\phi,s)$ are entire unless $\phi(\cdot)=|\cdot|_{{\mathbb{I}}_{F}}^{t}$ for some $t\in{\mathbb{C}}$; in the latter case, $\Lambda(\phi,s)$ has first-order poles precisely in s=ts=-t and s=1ts=1-t, and $L(\phi,s)$ has a unique first-order pole in s=1ts=1-t. Finally, there exists a holomorphic function $\epsilon(s)$ such that $\Lambda(\phi,s)$ satisfies the functional equation
@@
Using (25), the above functional equation can be rewritten as
(27) $\Lambda(\phi,s)=\epsilon(s)\Lambda(\bar{\phi},w(\phi)+1-s)\,.$
Remark 3.2.
For a Hecke character $\phi$ of $F$, define the Hecke character $\phi^{\perp}$ by putting
(28) $\phi^{\perp}(x):=(\phi(\alpha(x)))^{-1},\ \ \ x\in{\mathbb{I}}_{F}\,.$
If $\phi^{\perp}=\phi|\cdot|^{w(\phi)}_{{\mathbb{I}}_{F}}$, then
  1. (1)
    we have $\Lambda(\phi,s)=\Lambda(\bar{\phi},s)$, and the functional equation becomes
    $\Lambda(\phi,s)=\epsilon(\phi)\Lambda(\phi,w(\phi)+1-s)\$
  2. (2)
    $\epsilon(\frac{w+1}{2})$ is equal to $1$ or 1-1; we call this value the sign of the functional equation of $\phi$ and denote it by $\epsilon(\phi)$.
Now suppose that $w:=w(\phi)$ is an odd integer. If $\phi^{\perp}=\phi|\cdot|^{w}_{{\mathbb{I}}_{F}}$ and $\epsilon(\phi)=-1$, then an analysis of the $\Gamma$-factors shows that $L(\phi,s)$ has to vanish at the central point $s=\frac{w+1}{2}$.
@@
3.2. Motives
In order to discuss motives attached to Hecke characters, let us quickly review the notions that we will use about motives. For more details, the reader can consult [3] and [31].
Given a smooth, projective variety $X$ over a field $E$, denote by $\mathop{\rm CH}\nolimits^{n}(X)$ the Chow group of codimension $n$ algebraic cycles on $X$ modulo rational equivalence, with coefficients in $\mathbb{Q}$}. If $\dim_{X}$ is the dimension of $X$, then $\mathop{\rm CH}\nolimits^{\dim_{X}}(X\times X)$ becomes a ring under a composition law defined using intersection of cycles. Its elements are called (classes of) degree 0 self-correspondences on $X$, where, more generally, a correspondence between smooth projective varieties $X$ and $Y$ is an algebraic cycle on $X\times Y$.
The category $\mathop{\rm CHM}\nolimits(E)$ of Chow motives with rational coefficients over $E$ has objects of the form $(X,p,m)$ where $X$ is a smooth projective variety over $E$, $p$ is an idempotent element of $\mathop{\rm CH}\nolimits^{\dim_{X}}(X\times X)$, and $m$ is a integer. We often write $(X,p)$ for $(X,p,0)$. Morphisms are defined using rational equivalence classes of correspondences between smooth projective varieties, with maps $f:X\rightarrow Y$ between smooth projective varieties being interpreted as correspondences by identifying them with the transpose of their graph $\Gamma_{f}$. The category $\mathop{\rm CHM}\nolimits(E)$ is tensor and pseudo-abelian. Any Weil cohomology theory $H^{\bullet}(-)$ on the category $SmProj(E)$ of smooth projective varieties, towards graded vector spaces over a field $\mathbb{K}$ of characteristic zero, factors uniquely through a canonical functor
$SmProj(E)\rightarrow\mathop{\rm CHM}\nolimits(E)$
associating to a variety $X$ its Chow motive $M(X):=(X,\Gamma_{\mathop{\rm\tiny{id}}\nolimits_{X}})$. If there exists an embedding $\sigma$ of $E$ into $\mathbb{C}$, then one can take $\mathbb{K}=\mathbb{Q}$ and choose $H^{\bullet}(-)$ as the functor associating to $X$ the singular cohomology of the analytification of $X$. Then, the resulting functor from $\mathop{\rm CHM}\nolimits(E)$ to graded $\mathbb{Q}$}-vector spaces is called Betti realization and factors through the category of graded, pure $\mathbb{Q}$-Hodge structures. It sends $(X,p,m)$ to $pH^{\bullet}(X)(m)$. Here, $pH^{\bullet}(X)$ denotes the direct summand of $H^{\bullet}(X)$ obtained as the image of the idempotent correspondence $p$ acting on $H^{\bullet}(X)$, and ()(m)(-)(m) denotes the $m$th Tate twist in the Hodge-theoretical sense. If $E$ is of characteristic zero, then there exists a de Rham realization with coefficients in $E$. In general, one can fix a prime $\ell$ different from the characteristic of $E$, take $\mathbb{K}={\mathbb{Q}}_{\ell}$ and choose $H^{\bullet}(-)$ as being $\ell$-adic étale cohomology. The resulting $\ell$-adic realization $M_{\ell}$ of a Chow motive $M$ is then endowed with the structure of a $\ell$-adic Galois representation, which allows one to define an $L$-function $L(M,s)$.
To deal with more general varieties than smooth and projective ones, one enlarges $\mathop{\rm CHM}\nolimits(E)$ to a symmetric monoidal, $\mathbb{Q}$}-linear triangulated category of constructible motives over $E$, denoted $\mathop{\rm DM}\nolimits_{c}(E)$, equipped with a fully faithful embedding
$\mathop{\rm CHM}\nolimits(E)\hookrightarrow\mathop{\rm DM}\nolimits_{c}(E)\,.$
The category $\mathop{\rm DM}\nolimits_{c}(E)$ was first constructed by Voevodsky (who called it the category of geometrical motives and denoted it by $\mathop{\rm DM}\nolimits_{gm}(E)$) starting from the complexes over the additive category of smooth varieties equipped with finite correspondences, with no equivalence relation imposed on them. Any variety $X$ over $E$, even if non-projective or singular, has an associated motive $M(X)$ in $\mathop{\rm DM}\nolimits_{c}(E)$, identified with its Chow motive when $X$ is smooth and projective. When $E$ embeds into $\mathbb{C}$}, the Betti realization extends to $\mathop{\rm DM}\nolimits_{c}(E)$ and factors through a Hodge realization functor $R_{H}$ towards the category $\mathop{\rm MHS}\nolimits_{{\mathbb{Q}}}$ of mixed $\mathbb{Q}$}-Hodge structures.
For any field $\mathbb{K}$}, there are $\mathbb{K}$}-linear analogues $\mathop{\rm CHM}\nolimits(E)_{{\mathbb{K}}}$, $\mathop{\rm DM}\nolimits_{c}(E)_{{\mathbb{K}}}$, called with coefficients in $\mathbb{K}$}, of the above categories, with a Hodge realization $R_{H}$ factoring through $\mathop{\rm MHS}\nolimits_{{\mathbb{K}}}$. An object $A$ of $\mathop{\rm MHS}\nolimits_{\overline{{\mathbb{Q}}}}$ is said to be of geometric origin if $A$ belongs to the full Tannakian subcategory of $\mathop{\rm MHS}\nolimits_{\overline{{\mathbb{Q}}}}$ generated by the cohomology spaces of those objects which are in the image of $R_{H}$.
@@ 3.3. Motives for Hecke characters
By [17, Prop. 3.5], there exists a Chow motive $M_{\phi}$ with coefficients in a finite extension of $F^{\prime}$, whose realizations are the Hodge structure $H_{\phi}$ and the compatible system $(M_{\phi,\ell})_{\ell}$ (the construction runs as in [39, Ch.1, Thm. 4.1], where one obtains an object in the weaker category of absolute Hodge motives). Then, one has $L(\phi,s)=L(M_{\phi},s)$. To describe this motive, recall that by [17], for every abelian variety $A$, there exists a canonical Chow motive $h^{i}(A)$ realising the $i$-th cohomology of $A$ in any Weil cohomology theory. Then, the motives $M_{\phi}$ have the following form: if the weight $w$ of $\phi$ is positive, there exist an abelian variety $A_{\phi}$ over $F$, with complex multiplication by a commutative semisimple algebra $T$ over $F$, and an idempotent element $e_{\phi}$ of $T^{\otimes w}$, such that
$M_{\phi}=(h^{1}(A_{\phi})^{\otimes w},e_{\phi})\,.$
For any embedding $F\hookrightarrow{\mathbb{C}}$, we can consider the singular cohomology spaces $H^{\bullet}(A_{\phi})$, with coefficients in a suitable number field, of the analytification of $A_{\phi}$ with respect to the chosen embedding. The space $H^{1}(A_{\phi})^{\otimes w}$ has a canonical pure Hodge structure of weight $w$, with coefficients in the same number field. The Hodge structure $H_{\phi}$ can then be described as
(29) $H_{\phi}=e_{\phi}H^{1}(A_{\phi})^{\otimes w}\,.$
By the Künneth theorem, $H^{1}(A_{\phi})^{\otimes w}$ is a direct factor of $H^{w}(A_{\phi}^{w})$. The results of [17] imply that there exists an idempotent correspondence $e_{1,w}$ cutting out a direct factor of $h^{w}(A_{\phi}^{w})$ realizing to $H^{1}(A_{\phi})^{\otimes w}$, so that we can also write
$M_{\phi}=(h^{w}(A^{w}_{\phi}),e_{\phi}e_{1,w})$
and the Hodge structure $H_{\phi}$ can also be described as
$H_{\phi}=e_{\phi}e_{1,w}H^{w}(A^{w}_{\phi})\,.$